Two angles are complements and one lacks 8 degrees of being twice the other. What are the angles?
I tried: 90 = 2x - 8 + y
90 + 8 = 2x - 8 + 8 + y
98 = 2x + y
49 = x + y
Stuck here

Answer by checkley77(12569)

x+y=90 or x=90-y
x-8=2y
x=2y+8
2y+8=90-y
2y+y=90-8
3y=82
y=82/3
y=27 1/3 ans.
x=90-27 1/3=62 2/3 ans.
Proof:
62 2/3-8=2*27 1/3
54 2/3=54 2/3